# A deep learning approach for predicting multiple observables in Au+Au collisions at RHIC

### v2

@@ II Model framework
where $z_{n},x_{n},y_{n}$ are the coordinates of the $n$th neuron on the sphere centered at the coordinate origin, $R$ is the radius of the sphere, $N$ is the total number of neurons in the layer, $\phi$ is the golden ratio. With this distribution, we can determine the distance between neurons and implement the local connection between the QGP, chemical freeze-out and kinetic freeze-out layers. Each neuron in the chemical freeze-out layer is connected only to its seven nearest neighbors in the QGP layer, and the same principle is applied to the connection between the chemical freeze-out layer and kinetic freeze-out layer. This induces a structured sparse connectivity pattern that regularizes the network by suppressing spurious long-range couplings while preserving local correlations, playing a similar role to Dropout but in a more physically interpretable way.
@@ IV Summary
In this work we developed a physics-informed deep neural network to predict multiple bulk observables in Au+Au collisions at RHIC. The architecture is inspired by the space–time stages of a heavy-ion collision and incorporates locally connected hidden layers between QGP, chemical and kinetic freeze-out representations, together with a structured dual-input design encoding both initial collision characteristics and final-state kinematics. The network is trained exclusively on experimental measurements of charged-particle pseudorapidity density distributions, transverse-momentum spectra and elliptic flow over a broad range of collision energies and centralities.
@@ Acknowledgements.
We thank for the helpful discussions with Prof. Wei Dai, Prof. Ze-Fang Jiang and Prof. Dirk H. Rischke. The work is supported by the China Scholarship Council (CSC) (Jun-Qi Tao), the CUHK-Shenzhen University development fund under grant No. UDF01003041 and UDF03003041, the Shenzhen Peacock fund under No. 2023TC0179 and the Ministry of Science and the Technology of China under Grant No. 2024YFA1611004 (Kai Zhou), the National Natural Science Foundation of China (NSFC) under Grants No. 12535010 (Ben-Wei Zhang).

### v3

@@ II Model framework
where $z_{n},x_{n},y_{n}$ are the coordinates of the $n$th neuron on the sphere centered at the coordinate origin, $R$ is the radius of the sphere, $N$ is the total number of neurons in the layer, $\phi$ is the golden ratio. With this distribution, we can determine the distance between neurons and implement the local connection between the QGP, chemical freeze-out and kinetic freeze-out layers. Each neuron in the chemical freeze-out layer is connected only to its seven nearest neighbors in the QGP layer, and the same principle is applied to the connection between the chemical freeze-out layer and kinetic freeze-out layer. This induces a structured, sparse connectivity pattern that regularizes the network by suppressing spurious long-range couplings while preserving local correlations, playing a similar role to Dropout but in a more physically interpretable way. For the connection between other layers, the full connection is used.
@@ IV Summary
In this work, we developed a physics-informed deep neural network to predict multiple bulk observables in Au+Au collisions at RHIC. The architecture is inspired by the space–time stages of a heavy-ion collision and incorporates locally connected hidden layers between QGP, chemical and kinetic freeze-out representations, together with a structured dual-input design encoding both initial collision characteristics and final-state kinematics. The network is trained exclusively on experimental measurements of charged-particle pseudorapidity density distributions, transverse-momentum spectra and elliptic flow over a broad range of collision energies and centralities.
@@ Acknowledgements.
We thank Prof. Wei Dai, Prof. Ze-Fang Jiang and Prof. Dirk H. Rischke for helpful discussions. This work is supported by the China Scholarship Council (CSC) under Grant No. 202406770003 (Jun-Qi Tao), the CUHK-Shenzhen University Development Fund under Grant Nos. UDF01003041 and UDF03003041, the Shenzhen Peacock Fund under Grant No. 2023TC0179, the Ministry of Science and Technology of China under Grant No. 2024YFA1611004 (Kai Zhou) and the National Natural Science Foundation of China (NSFC) under Grant No. 12535010 (Ben-Wei Zhang).